OpenStudy (anonymous):

Can someone help teach me how to solve this? (give me a second to make the rational expression)

OpenStudy (anonymous):

1. Simplify the expression and state the restrictions. \[\left( \frac{ 9x^7 }{ 15x^5 } \right)\left(\frac{ 5x^2 }{ 45x^3 } \right)\]

OpenStudy (anonymous):

I DON"T just want the answer. I really need someone to teach me the steps. @Mertsj

OpenStudy (mertsj):

1. The denominator cannot be 0 so exclude the value of x that would make the denominator 0. 2. Reduce each fraction by cancelling the factors that are common to numerator and denominator 3. Multiply the remaining two fractions.

OpenStudy (anonymous):

okay. the lesson says factor first, so after factoring i got \[\left( \frac{ 3x^2 }{ 5 } \right)\left( \frac{ 1 }{ 9x } \right)\] then it says after I factor i can find the restrictions but the example doesn't look like my equation so im stuck. the example is

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

You should tell the restrictions BEFORE you cancel the common factors.

OpenStudy (mertsj):

In your case, x cannot be 0 because then the denominator would be 0 and you would be arrested and sent to the gulag.

OpenStudy (mertsj):

So now you need to cancel the 3 and the x

OpenStudy (mertsj):

What you have done so far is correct.

OpenStudy (anonymous):

how do I cancel the three and the x

OpenStudy (mertsj):